# A QUBO Formulation for the Generalized LinkedIn Queens and Takuzu/Tango Game

### v2

@@ 2.1 Problem definition
  • Diagonal condition: there cannot be two queens in the same diagonal (Fig. 1 c):
    $\sum_{(k,l)\in D_{i,j}}x_{kl}\leq 1\quad\forall i,j\in[1,N],$
    being $D_{ij}$ the set of cells on the diagonals of cell $(i,j)$, including it.
@@
  • Diagonal condition: there cannot be more than one queen in the same adjacent diagonal (Fig. 1 d): $x_{ij}+X_{i+1,j+1}+X_{i-1,j+1}+X_{i+1,j-1}+X_{i-1,j-1}\leq 1\quad\forall i,j\in[1,N]$, being $X_{kl}=x_{kl}$ if $k,l\in[1,N]$ and 00 otherwise.
@@
An additional condition in this case is that there are a number $N_{I}$ of queens already placed in cells $A_{I}$, as an initial condition. This can be addressed by forcing the variables associated with these cells to be forcibly 1, which leads to all other adjacent variables in the same row, in the same column and in the same region to be forcibly 0. This translates to
@@ 2.2 QUBO resolution
For the restriction of the diagonals we use a term of the type
$Q_{d}=\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl},$ (3)
so that, for each occupied cell $(i,j)$, if another cell on its diagonal is occupied, this term sums to 1, otherwise it sums to 0. In this case we define $D^{\prime}_{i,j}$ as the cells that are on the same diagonals as cell $(i,j)$ such that their row is greater than $i$. In this way we avoid adding twice over the same pair of cells and we also avoid the term $x_{ij}^{2}=x_{ij}$, which would give us 1 whenever that cell was occupied. As only cells in rows below the cell where there is a queen are computed, we do not need to include the condition in the last row for $i$.
@@
$Q=$ $Q_{r}+Q_{c}+Q_{d}=$
$=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}.$ (4)
@@
For the diagonal condition in this case, we have the same term, redefining in this case $D^{\prime}_{ij}$ as the diagonal cells adjacent to cell $(i,j)$, such that their row is greater than $i$, instead of the whole diagonals, and which have not already been fixed by the initial condition. Because of the initial condition, we have to construct the term as
@@
In this case we only consider cells that are not fixed.
@@
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=$
$=$ $\sum_{i\in A_{r}}\left(1-\sum_{j\in A_{c}(i)}x_{ij}\right)^{2}+$
$+$ $\sum_{j\in A_{c}}\left(1-\sum_{i\in A_{r}(j)}x_{ij}\right)^{2}+$
$+$ $\sum_{(i,j)\in A^{\prime}_{C}}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k\in A_{R}}\left(1-\sum_{(i,j)\in R_{k}|(i,j)\in A_{C}}x_{ij}\right)^{2}.$ (9)
@@ 2.3.2 Distant diagonals constraint
The term that implements this is similar to the ones shown above in (3), as $d_{ij}$ only change the set $D_{i,j}$. In this case, because of the non-symmetry of the required diagonal distance between cells, $D^{\prime\prime}$ has to be used in the summation, instead of $D^{\prime}$. $D^{\prime\prime}_{ij}$ contains all the cells of $D_{ij}$ but $(i,j)$.
$Q_{d}=\sum_{i,j}^{N,M}\sum_{(k,l)\in D^{\prime\prime}_{i,j}}x_{ij}x_{kl},$ (12)
Given this abstract formulation, this condition can be generalized so that the 4 diagonals do not have the same distance, that instead of being continuous lines they can be skipped lines, that instead of being 4 diagonals they are $n_{d}$ lines with different slopes, or that they are directly abstract regions, since the only thing that change is the definition of the $D_{ij}$ over which to sum for each cell.
@@ 2.3.3 Number of queens per region
$Q$ $=Q_{g}+Q_{d}=\sum_{k=1}^{N_{R}+N+M}\left(q_{k}-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N,M}\sum_{k,l\in D_{i,j}}x_{ij}x_{kl}$ (14)
where the set of regions $R^{\prime}$ is contains the $N_{R}$ regions $R_{k}$, the $N$ rows and the $M$ columns.
@@ 2.3.4 Two possible numbers of queens per region
As can be seen, the diagonal constraints are only particular cases of this constraint in which there is one region for each cell, which would be its diagonals. The value $v_{k}$ would be 0 for all the regions, since the number of queens must be 0 or 1. We can rewrite the eq (2.3.3), with $N_{R}$ regions $R_{k}$ with a single possible number of queens and $N_{V}$ regions $V_{k}$ with two consecutive possible numbers of queens, as
$Q=$ $Q_{g}+Q_{v}=\sum_{k=1}^{N_{R}}\left(q_{k}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}+$
$+$ $\sum_{k=1}^{N_{V}}\left(v_{k}+\frac{1}{2}-\sum_{(i,j)\in V_{k}}x_{ij}\right)^{2}.$ (16)
@@ 2.4 Definition of generalised N-queens
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed $N_{I}$ initial queens in positions given by the set $A_{I}$, and divided into $N_{R}$ regions, which can overlap each other, such that in $k$-th region there must be $q_{k}$ queens if $t_{k}=0$ and $q_{k}$ or $q_{k}+1$ queens if $t_{k}=1$, and we have initially placed $p_{k}$ queens, the objective will be to place all the possible queens on the board such that the given restrictions are satisfied.
@@
$Q=\sum_{k=1}^{N_{R}}\left(q_{k}-p_{k}+\frac{t_{k}}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2},$ (18)
where $R_{k}$ is the set of cells included in the $k$-th region that have not been set in the initial condition and that do not belong to a region where the number of queens available after the initial queens have been placed is 0.
@@ 2.4.1 N-queens with regions
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=$
$=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{R}}\left(1-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (19)
@@ 2.4.2 N-queens with soft regions
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{v}=$
$=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{R}}\left(\frac{1}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (20)
@@ 2.5 Tents & Trees
We can translate this problem to our formalism with the following considerations:
  • Each tree cell is no longer considered as an optimizing variable and is fixed to 0.
  • The $k$-th tree creates a region $R_{k}$, given by the 4 cells it has adjacent to it, such that in this region there must be at least 1 tent. In the cells not included by any of the regions, we fix the variables to 0.
  • The tents act as queens, but can only attack the 8 adjacent cells.
Our variables will be a set of binary $x_{ij}$ indicating that in row cell $i$ and column $j$ there is no tent if equal to 0 and there is if equal to 1.
The terms imposing the number of tents per row and column are exactly as we posed in the general case in (10) and (11). For the non-contact constraint between tents, we are going to use the term of (12), but making $D^{\prime}_{i,j}$, in addition to taking into account the adjacent diagonals, take into account the vertical and horizontal adjacencies. To reduce the number of terms, we will not take into account the cells above the tent, because if a tent has another one above it, the one above it will see it below it.
For the regions given by the trees, we will have to use a term of type (15) with $v_{k}=1$. This is because if there were 3 or 4 tents surrounding the tree, there would necessarily be 2 of them in diagonal contact. Therefore, there can only be 1 or 2 tents in that region.
Thus, the QUBO of the problem is
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=$
$=$ $\sum_{i=1}^{N}\left(N_{r,i}-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(N_{c,j}-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{t}}\left(\frac{3}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (22)
As in the previous cases, the minimum energy will be of $\frac{N_{t}}{4}$.

### v3

@@ 1 Introduction
1.1 QUBO preliminaries
For binary variables $x_{i}\in\{0,1\}$ we will repeatedly use the identity $x_{i}^{2}=x_{i}$, so every expanded penalty below is a genuine QUBO with at most linear and quadratic terms. Let
$s=\sum_{i\in I}x_{i}$
and let $A>0$ be a penalty coefficient.
  • Exact constraint $s=k$:
    $P_{=k}(s)=A(s-k)^{2}.$ (1)
    Since $s$ is an integer, $P_{=k}(s)\geq 0$ and it vanishes exactly when $s=k$.
  • At most one active variable:
    $P_{\leq 1}(I)=A\sum_{\begin{subarray}{c}i<j\\ i,j\in I\end{subarray}}x_{i}x_{j}.$ (2)
    Each product contributes 1 exactly when two variables are simultaneously active, so the minimizers are precisely the assignments with $\sum_{i\in I}x_{i}\leq 1$.
  • Exactly $q$ or $q+1$ active variables, without slack variables:
    $P_{q,q+1}(s)=A\left(s-q-\frac{1}{2}\right)^{2}.$ (3)
    Because $s$ is integer, the minimizers of (3) are exactly $s\in\{q,q+1\}$. Its minimum value is $A/4$, not 0. If one wants zero energy on feasible assignments, one may subtract the constant $A/4$ or equivalently use
    $P^{\prime}_{q,q+1}(s)=A(s-q)(s-q-1),$ (4)
    which has the same minimizers and differs from (3) only by the additive constant A/4-A/4.
  • No three equal consecutive values. For binary variables $a,b,c$ let $s=a+b+c$:
    $P_{\mathrm{triple}}(a,b,c)=A(s-1)(s-2).$ (5)
    Since $s\in\{0,1,2,3\}$, this penalty is positive only for $s=0$ and $s=3$, namely for the forbidden patterns $000$ and $111$, and it is zero for the other six assignments.
  • Equality $x=y$:
    $P_{=}(x,y)=A(x-y)^{2}.$ (6)
    This vanishes exactly when $x=y$.
  • Difference $x\neq y$:
    $P_{\neq}(x,y)=A(x+y-1)^{2}.$ (7)
    For binary variables, x+y1=0x+y-1=0 if and only if $(x,y)\in\{(0,1),(1,0)\}$.
  • Conflict graph. If $E$ is a set of forbidden simultaneous activations,
    $P_{E}(x)=A\sum_{\{u,v\}\in E}x_{u}x_{v}.$ (8)
    Each conflicting edge contributes exactly when both endpoints are active, so the minimizers are the independent sets of the conflict graph.
These standard penalties will be reused throughout the paper, sometimes up to additive constants, which do not affect the minimizing assignments.
@@ 2.1 Problem definition
  • Diagonal condition: there cannot be two queens on the same attacking diagonal (Fig. 1 c), namely
    $x_{ij}+x_{kl}\leq 1\quad\forall(i,j)\neq(k,l)\ \text{with}\ |i-k|=|j-l|.$
    We will use $D_{ij}$ only to denote the cells diagonally aligned with $(i,j)$, as an indexing device for these pairwise conflicts, not as a single region on which one imposes an at-most-one condition over the union of both diagonals through $(i,j)$.
@@
  • Diagonal condition: there cannot be more than one queen in the same adjacent diagonal (Fig. 1 d): $x_{ij}+X_{i+1,j+1}+X_{i-1,j+1}+X_{i+1,j-1}+X_{i-1,j-1}\leq 1\quad\forall i,j\in[1,N]$, being $X_{kl}=x_{kl}$ if $k,l\in[1,N]$ and 0 otherwise.
@@
An additional condition in this case is that there are a number $N_{I}$ of queens already placed in cells $A_{I}$, as an initial condition. This can be addressed by forcing the variables associated with these cells to be forcibly 1, which leads to all other conflicting variables in the same row, in the same column, in the same region and in the adjacent diagonals to be forcibly 0. This translates to
@@
    • Adjacent diagonals: $x_{ij}=0$ for every cell $(i,j)$ with |il|=|jm|=1|i-l|=|j-m|=1 and $\forall(l,m)\in A_{I}$.
@@ 2.2 QUBO resolution
For the restriction of the diagonals we use a conflict-graph term of the type
$Q_{d}=\sum_{\{u,v\}\in E_{\mathrm{diag}}}x_{u}x_{v}=\sum_{i,j=1}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl},$ (11)
where $E_{\mathrm{diag}}$ is the set of unordered pairs of cells lying on a common attacking diagonal. If we keep the notation $D^{\prime}_{i,j}$, it must be understood as the set of cells below $(i,j)$ that are in diagonal conflict with it, so that the sum counts conflicting pairs exactly once. In particular, we are not imposing an at-most-one condition on an arbitrary union of the two diagonals through a cell, which would be stronger than the actual N-queens rule.
@@
$Q=Q_{r}+Q_{c}+Q_{d}=\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+\sum_{i,j=1}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}.$ (12)
@@
For the diagonal condition in this case, we again use a pairwise conflict term, now restricted to adjacent diagonals. Thus $D^{\prime}_{ij}$ contains only the diagonally adjacent cells in rows below $i$ that remain unfixed after preprocessing. Because of the initial condition, we have to construct the term as
@@
so that every remaining adjacent-diagonal conflict is still counted exactly once. In this case we only consider cells that are not fixed.
@@
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=\sum_{i\in A_{r}}\left(1-\sum_{j\in A_{c}(i)}x_{ij}\right)^{2}+\sum_{j\in A_{c}}\left(1-\sum_{i\in A_{r}(j)}x_{ij}\right)^{2}+$
$+$ $\sum_{(i,j)\in A^{\prime}_{C}}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+\sum_{k\in A_{R}}\left(1-\sum_{(i,j)\in R_{k}|(i,j)\in A_{C}}x_{ij}\right)^{2}.$ (17)
@@
Proposition 2.1.
After preprocessing the pre-placed queens by fixing $x_{lm}=1$ for $(l,m)\in A_{I}$ and fixing to 0 every remaining cell in the same row, column, region or adjacent diagonal, the minimum-energy states of the reduced LQueens QUBO coincide with the solutions of the original instance restricted to the non-fixed variables.
Proof.
Every constraint involving a pre-placed queen is already satisfied by the preprocessing, because every cell that could violate such a constraint has been fixed to 0. Hence all removed terms are constant on the reduced instance. Conversely, any assignment of the remaining variables violates the original puzzle if and only if it violates one of the surviving row, column, adjacent-diagonal or region constraints, which are precisely the terms kept in the reduced QUBO. ∎
@@ 2.3.2 Distant diagonals constraint
The term that implements this remains a conflict-graph term. Let $E_{d}$ be the set of unordered pairs of cells such that one of them lies on one of the allowed diagonals of the other and at the required distance. Then the penalty is obtained by summing one product per conflicting pair. If one prefers to keep a set notation, $D^{\prime\prime}_{ij}$ should only enumerate the cells that conflict with $(i,j)$, not define a single region with an at-most-one condition over a union of diagonals.
$Q_{d}=\sum_{\{u,v\}\in E_{d}}x_{u}x_{v}.$ (20)
Given this abstract formulation, this condition can be generalized so that the 4 diagonals do not have the same distance, that instead of being continuous lines they can be skipped lines, or that instead of being 4 diagonals they are $n_{d}$ attack lines with different slopes. In every case, what matters is the underlying conflict graph, not an arbitrary region obtained by merging unrelated lines.
@@ 2.3.3 Number of queens per region
$Q=Q_{g}+Q_{d}=\sum_{k=1}^{N_{R}+N+M}\left(q_{k}-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}+\sum_{\{u,v\}\in E_{d}}x_{u}x_{v},$ (22)
where the set of regions $R^{\prime}$ contains the $N_{R}$ regions $R_{k}$, the $N$ rows and the $M$ columns, while $E_{d}$ is the conflict graph describing the diagonal-style attacks.
@@ 2.3.4 Two possible numbers of queens per region
The term (23) is useful for genuine region constraints with two consecutive admissible occupancies. Diagonal attack constraints are instead better expressed through pairwise conflict graphs as in (11) and (20). We can rewrite the eq (22), with $N_{R}$ regions $R_{k}$ with a single possible number of queens and $N_{V}$ regions $V_{k}$ with two consecutive possible numbers of queens, as
$Q=Q_{g}+Q_{v}=\sum_{k=1}^{N_{R}}\left(q_{k}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}+\sum_{k=1}^{N_{V}}\left(v_{k}+\frac{1}{2}-\sum_{(i,j)\in V_{k}}x_{ij}\right)^{2}.$ (24)
@@
This compact expression summarizes only the region part of the objective. Whenever diagonal or line-of-sight attacks are present, the corresponding conflict-graph penalty must still be added separately. Moreover, for the terms with $t_{k}=1$, feasible assignments share a constant offset of $1/4$ per such region unless one uses the shifted form from (4).
@@ 2.4 Definition of generalised N-queens
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed $N_{I}$ initial queens in positions given by the set $A_{I}$, and divided into $N_{R}$ regions, which can overlap each other, such that in the $k$-th region there must be $q_{k}$ queens if $t_{k}=0$ and $q_{k}$ or $q_{k}+1$ queens if $t_{k}=1$, and we have initially placed $p_{k}$ queens, the objective will be to place all the possible queens on the board such that the given region restrictions and pairwise attack restrictions are satisfied.
@@
$Q=\sum_{k=1}^{N_{R}}\left(q_{k}-p_{k}+\frac{t_{k}}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}+\sum_{\{u,v\}\in E_{d}}x_{u}x_{v},$ (26)
where $R_{k}$ is the set of cells included in the $k$-th region that have not been set in the initial condition and that do not belong to a region where the number of queens available after the initial queens have been placed is 0, and $E_{d}$ is the chosen conflict graph for the diagonal or line-of-sight restrictions. If $t_{k}=1$, the corresponding feasible region contributes a constant offset of $1/4$ unless one uses the shifted form from (4).
@@ 2.4.1 N-queens with regions
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+\sum_{k=1}^{N_{R}}\left(1-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (27)
@@ 2.4.2 N-queens with soft regions
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{v}=\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+\sum_{k=1}^{N_{R}}\left(\frac{1}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (28)
Here the feasible soft-region assignments share a constant offset of $N_{R}/4$ because of the fractional terms.
@@ 2.5 Tents & Trees
If one directly adapts the regional formalism, each tree neighborhood can be treated as a region that should contain one or two tents, and the no-touch condition can be enforced through pairwise conflicts between king-adjacent cells. This gives a useful simplified relaxation, but it is not exact in general because overlapping neighborhoods do not enforce a one-to-one assignment between trees and tents.
For an exact formulation, let $F$ be the set of free cells, excluding the tree cells themselves. For each $c\in F$, introduce a binary variable $x_{c}$ indicating whether there is a tent at cell $c$. For each tree $t$ and each free cell $c$ orthogonally adjacent to $t$, introduce a binary assignment variable $y_{t,c}$, and denote by $N(t)\subseteq F$ the set of free cells orthogonally adjacent to $t$. Let $E_{\mathrm{king}}$ be the conflict graph on $F$ connecting two cells when they are adjacent orthogonally or diagonally.
The exact penalties are then
$Q_{\mathrm{tree}}=A\sum_{t}\left(\sum_{c\in N(t)}y_{t,c}-1\right)^{2},$ (30)
$Q_{\mathrm{assign}}=A\sum_{c\in F}\left(x_{c}-\sum_{t:\,c\in N(t)}y_{t,c}\right)^{2},$ (31)
$Q_{\mathrm{touch}}=A\sum_{\{c,c^{\prime}\}\in E_{\mathrm{king}}}x_{c}x_{c^{\prime}},$ (32)
$Q_{\mathrm{row}}=A\sum_{i=1}^{N}\left(\sum_{j:\,(i,j)\in F}x_{ij}-N_{r,i}\right)^{2},$ (33)
$Q_{\mathrm{col}}=A\sum_{j=1}^{N}\left(\sum_{i:\,(i,j)\in F}x_{ij}-N_{c,j}\right)^{2}.$ (34)
If a particular puzzle does not specify row or column counts, the corresponding term can simply be omitted.
Therefore, an exact QUBO for Tents & Trees is
$Q_{Tents}^{\mathrm{exact}}=Q_{\mathrm{tree}}+Q_{\mathrm{assign}}+Q_{\mathrm{touch}}+Q_{\mathrm{row}}+Q_{\mathrm{col}}.$ (35)
Proposition 2.2.
The zero-energy states of $Q_{Tents}^{\mathrm{exact}}$ are exactly the standard solutions of Tents & Trees.
Proof.
$Q_{\mathrm{tree}}$ gives exactly one assigned tent to each tree. $Q_{\mathrm{assign}}$ guarantees that every occupied tent cell is assigned to exactly one adjacent tree and that no assignment is made to a cell without a tent. $Q_{\mathrm{touch}}$ forbids orthogonal and diagonal contact between tents, while $Q_{\mathrm{row}}$ and $Q_{\mathrm{col}}$ impose the row and column counts. Conversely, every valid puzzle solution induces a unique compatible assignment of each tent to its adjacent tree, so all five penalties vanish. ∎